### v1

# Monotonic Paraphrasing Improves Generalization of Language Model Prompting

Abstract
Performance of large language models (LLMs) may vary with different prompts or instructions of even the same task. One commonly recognized factor for this phenomenon is the model’s familiarity with the given prompt or instruction, which is typically estimated by its perplexity. However, finding the prompt with the lowest perplexity is challenging, given the enormous space of possible prompting phrases. In this paper, we propose monotonic paraphrasing (MonoPara), an end-to-end decoding strategy that paraphrases given prompts or instructions into their lower perplexity counterparts based on an ensemble of a paraphrase LM for prompt (or instruction) rewriting, and a target LM (i.e. the prompt or instruction executor) that constrains the generation for lower perplexity. The ensemble decoding process can efficiently paraphrase the original prompt without altering its semantic meaning, while monotonically decrease the perplexity of each generation as calculated by the target LM. We explore in detail both greedy and search-based decoding as two alternative decoding schemes of MonoPara. Notably, MonoPara does not require any training and can monotonically lower the perplexity of the paraphrased prompt or instruction, leading to improved performance of zero-shot LM prompting as evaluated on a wide selection of tasks. In addition, MonoPara is also shown to effectively improve LMs’ generalization on perturbed and unseen task instructions.
1 Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in zero-shot decision making Gonen et al. (2023); Schick and Schütze (2021); Brown et al. (2020) and instruction following Jiang et al. (2023); Köpf et al. (2023); Touvron et al. (2023b); Taori et al. (2023); Chiang et al. (2023); Ouyang et al. (2022). However, there can be significant variance in the performance of seemingly similar prompts Zhao et al. (2021); Lu et al. (2022c); Webson and Pavlick (2022); Gonen et al. (2023). Despite efforts of studies on prompting LMs Shin et al. (2020); Li and Liang (2021); Gao et al. (2021); Ding et al. (2022); Sanh et al. (2021); Kojima et al. (2022), it is still challenging to develop high-quality prompts that can induce better performance for varying tasks on evolving models in an effort-saving manner.
One consensus reached by recent studies is the inverse relationship between a prompt’s perplexity and its task performance Kumar et al. (2023); Gonen et al. (2023). This stems from the intuition that the frequency of a prompt (or an instruction) [^1] We use the terms “prompt” and “instruction” interchangeably in this paper as they both refer to a natural language command for zero-shot inference. appearing in the (pre-)training data positively influences the model’s familiarity with it, thereby enhancing its capability to perform the described task Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023). As a result, existing attempts use specific criteria, especially perplexity Gonen et al. (2023); Ye and Durrett (2023); Kumar et al. (2023), for selecting prompts from a collection of candidates. For example, Gonen et al. (2023) builds a large prompt pool for each task and selects the one with the lowest perplexity. However, the performance of such a method is limited by the size, quality, and even availability of the candidate pool. Searching the prompt with the lowest perplexity for a particular task remains challenging due to the vast expanse of potential prompts.
To address this challenge, we propose a novel progressive prompt refining approach, namely monotonic paraphrasing (MonoPara), that can proactively rewrite the given prompt into its low-perplexity counterpart without compromising its expressiveness of the task. During prompt (or instruction) refinement, MonoPara conducts an ensemble decoding process of a paraphrasing model together with the target LM. The paraphrase model is instructed to rewrite the given prompt of a specific task, while the target model, which is later on used for task inference, provides a constraint that aims to lower the perplexity of the generation. MonoPara iteratively decodes tokens based on the ensemble of these two models. Intuitively, the paraphrase model can genuinely paraphrase the original prompt without altering its semantic meaning and remain instructive for the given task, while the target model restricts the search space of paraphrased tokens to those with low perplexity, resulting in a task prompt that the target model is more familiar with.
Further, for the decoding process of MonoPara, we explore two decoding schemes. In addition to greedy decoding (Section 2.3) that iteratively generates the next token based on the weighted combination of predicted probabilities from both models, we also explored search-based decoding (Section 2.4). The search-based decoding scheme follows the look-ahead decoding paradigm Lu et al. (2022b), which keeps several sequence candidates and maintains the one with the lowest perplexity scored by the target model. Compared to recent prompt refinement approaches Gonen et al. (2023); Kumar et al. (2023); Shum et al. (2023), MonoPara needs zero training effort and far less computational cost, and do not require pre-existence of multiple prompt or instruction candidates Gonen et al. (2023). We find that by an ensemble of two LMs, we are able to consistently decode lower perplexity prompts, which would in turn bring about better generalization of LMs on downstream tasks.
Our contributions are three-fold. First, we raise the intuition that the greedy generation of prompted paraphrasing is not monotonic in perplexity (Section 2.1). Second, we explore and compare two distinct decoding strategies that proactively refine prompts by monotonically lowering their perplexity and maintaining their expressiveness of the task. Third, we conduct experimentation for both prompt refinement for regular LMs and instruction refinement for instruction-tuned LMs to illustrate that monotonic paraphrasing improves the generalization of LM prompting.
Figure 1: Perplexity of $x_{para}$ as the output paraphrase of $P_{para}$ vs. as the input prompt of $P_{tar}$ for the AG News dataset with Mistral 7B as both $P_{para}$ and $P_{tar}$. Each point stands for a different prompt $x_{para}$. A low-perplexity paraphrase does not necessarily result in a low-perplexity prompt for the target model.
Figure 2: Two explored decoding schemes of MonoPara. Ensemble-based decoding (bottom) combines the token probabilities from the paraphrase model and the target model in each decoding step. Search-based decoding (top) further leverages look-ahead decoding to consider the potential future impact of current choices.
2 Monotonic Paraphrasing
We propose MonoPara as an ensemble-based decoding method that can refine a given prompt by paraphrasing it into a low-perplexity counterpart. We first provide intuition of MonoPara in Section 2.1 and the basic of prompt paraphrasing in Section 2.2. Then, we formally define two decoding schemes of MonoPara in Section 2.3 and Section 2.4.
2.1 Intuition and Methodology
Assume that the user wants to address a task specified by a given prompt $x_{ori}$ using a target LM $P_{tar}$. MonoPara seeks to refine the given prompt to its low-perplexity counterpart $x_{para}$ (i.e. a more familiar prompt to $P_{tar}$), thereby enhancing model performance. However, simply paraphrasing the original prompt with a paraphrase model $P_{para}$ does not necessarily result in a low-perplexity counterpart. This discrepancy arises from a mismatch between the perplexity of the output $x_{para}$ from the paraphrase model and the perplexity observed when using $x_{para}$ as the input prompt for the target model. As shown in Figure 1, there is a lack of significant correlation between the perplexity of $x_{para}$ as output in $P_{para}$ and that of $x_{para}$ as input in $P_{tar}$, even if we use the same model as the paraphrase and target model ($P_{para}=P_{tar}$). This observation indicates that the minimal perplexity of a paraphrased prompt could not be achieved by the paraphrase model alone.
Based on the above fact, the goal of MonoPara is to take on the semantic hints from the paraphrase model while following the constraint of perplexity from the target model. Specifically, the paraphrase model generates paraphrases for a given prompt, while the target model constrains the paraphrase perplexity during the generation process. With an ensemble of these two models, we can leverage the strengths of both: the guided content generation capability of the paraphrase model and the low-perplexity prompt constraint of the target model. MonoPara is thus able to produce paraphrases with lower perplexity by synergizing the benefits of both models.
2.2 Prompt Paraphrasing
We first introduce prompt paraphrasing without perplexity constraint. We consider using an LLM for prompt paraphrasing in a zero-shot manner, where the model is instructed to paraphrase the original prompt $x_{ori}$ to a fluent and coherent counterpart $x_{para}$ with a system prompt $x_{sys}$, such as the following one:
Specifically, we denote the input of paraphrase model, including the concatenation of the system prompt $x_{sys}$ and the original prompt $x_{ori}$ that the model is instructed to rewrite, as $[x_{sys},x_{ori}]=(x_{1},\dots,x_{n})$, where $x_{i}$ is a token in the vocabulary $V$. Suppose the decoder of a pre-trained autoregressive language model $P_{para}$ generates continuations of length $m$ as the requested paraphrase, denoted as $x_{para}=(x_{n+1},\dots,x_{n+m})$, based on the system prompt $x_{sys}$ and original prompt $x_{ori}$. At decoding time, $P_{para}$ iteratively decodes one token at a time by conditioning on the preceding context:
$P_{para}(x_{para}|x_{sys},x_{ori})=\prod_{i=n+1}^{n+m}P_{para}(x_{i}|x_{<i}),$
where $P_{para}(x_{i}|x_{<i})$ is the predicted next token probability.
2.3 Ensemble-based (Greedy) MonoPara
To introduce perplexity constraint from the target model to the paraphrase model, here we introduce the ensemble-based greedy decoding for monotonic paraphrasing (as shown in the lower half of Figure 2). Based on the intuition delivered in Section 2.1, MonoPara decodes the paraphrase iteratively, and the next token is selected based on the combination of predicted probabilities from the two LMs:
$x_{next}$ $=\underset{x_{i}\in V}{\mathrm{argmax}}(\alpha\cdot\log P_{tar}(x_{i}|x_{gen})$
$+(1-\alpha)\cdot\log P_{para}(x_{i}|x_{sys},x_{ori},x_{gen})),$
where $x_{gen}$ denotes the sequence of generated tokens, and $\alpha$ determines the coefficient between the two models (Algorithm 1).
By combining the prediction probabilities of both models, the ensemble can navigate the complex landscape of language generation more effectively and precisely. The paraphrase model $P_{para}$ keeps the rewritten prompt (or instruction) remain with the original intent, while the target model $P_{tar}$ evaluates the paraphrase’s likelihood or confidence, up-weighting any generation samples that lead towards low-perplexity outcomes. Based on the ensemble, the next token is selected by maximizing the weighted sum of logarithmic probabilities from both models, considering both the instruction and the generated text thus far.
As a coefficient, $\alpha$ balances the contribution of each model, ensuring that the final output satisfies both semantic fidelity to the prompt intent and linguistic familiarity to the target model. Higher $\alpha$ increases the weight of the logarithmic probability derived from the target model $P_{tar}$ and cuts down on the semantic hints for $P_{para}$, which raises the risk of sacrificing the semantic fidelity of the generation. Conversely, a lower $\alpha$ shifts the emphasis towards $P_{para}$, placing greater value on ensuring that the generated text closely follows the instructional context and semantic intent from $P_{para}$. In this case, the decoding process lacks the constraint of perplexity from $P_{tar}$ and sacrifices the linguistic familiarity of the generated paraphrase for the target model.
Algorithm 1 Ensemble-based Decoding
1:
2:$x_{sys}$triangleright System prompt for paraphrase
3:$V$triangleright Vocabulary
4:$P_{tar},P_{para}$triangleright Target and paraphrase models
5:$\alpha$triangleright Weighting factor
6:
7:$x_{gen}$triangleright Generated text sequence
8:Initialize $x_{gen}\leftarrow\emptyset$
9:$x_{next}\leftarrow\underset{x_{i}\in V}{\mathrm{argmax}}\ P_{para}(x_{i}|x_{sys}\oplus x_{ori})$triangleright Generate first token with $P_{para}$
10:$x_{gen}\leftarrow x_{gen}\oplus x_{next}$
11:while not end of sequence do
12:     Compute weighted log probabilities for each $x_{i}\in V$:
13:     $score_{tar}\leftarrow\alpha\cdot\log P_{tar}(x_{i}|x_{gen})$
14:     $score_{para}\leftarrow(1-\alpha)\cdot\log P_{para}(x_{i}|x_{sys}\oplus x_{ori}\oplus x_{gen})$
15:     $x_{next}\leftarrow\underset{x_{i}\in V}{\mathrm{argmax}}\ (score_{tar}+score_{para})$
16:     Append $x_{next}$ to $x_{gen}$
17:end while
18:return $x_{gen}$
2.4 Search-based MonoPara
To further enhance the efficiency of decoding prompts of lower perplexity, we define the search-based decoding strategy for MonoPara (as shown in the upper half of Figure 2). To take more steps of token generation into account and expand the search space, we leverage look-ahead decoding to search for the sequence that exhibits the lowest perplexity from an even broader candidate space. Following Lu et al. (2022a), each step of our search-based decoding consists of (i) expanding a set of candidate next-tokens, (ii) scoring each candidate, and (iii) selecting the $k$ best candidates (Algorithm 2):
$X^{\prime}_{m}$ =$=\{\mathrm{x}_{<m}\circ x_{m}|\mathrm{x}_{<m}\in X_{m-1},x_{m}\in V\},$
$X_{m}$ $=\underset{(\mathrm{x}_{<m},x_{m})\in X^{\prime}_{m}}{\mathrm{\arg\,topk}}\{f(\mathrm{x}_{<m},x_{m})\},$
where $x_{m}$ is a candidate predicted by paraphrase model, and $f(\cdot)$ is a scoring function that returns the perplexity of $\mathrm{x}_{<m}\circ x_{m}$ evaluated by the target model:
$f(\mathrm{x}_{<m},x_{m})=\left(\prod_{i=1}^{m}\frac{1}{P_{tar}(x_{i}|x_{<i})}\right)^{\frac{1}{m}}.$ (1)
Specifically, at each step $m$ of the decoding process, the current set of sequences $X_{m-1}$ is expanded by appending the most probable next-token $x_{m}$, predicted by the paraphrase model $P_{para}$, from vocabulary $V$ to each sequence $\mathrm{x}_{<m}$. This results in a temporary expanded set of candidate sequences $X^{\prime}_{m}$. Each candidate sequence in $X^{\prime}_{m}$ is then evaluated by the scoring function $f(\cdot)$ shown in Equation 1, which calculates the perplexity of a sequence by the target model $P_{tar}$. Since a paraphrase of lower perplexity is preferred, the $k$ sequences with the best scores (i.e., lowest perplexity) are selected to form the set $X_{m}$, narrowing down the choices to the $k$ most promising sequences for continuation.
In comparison to greedy decoding, this look-ahead decoding mechanism evaluates a broader range of potential future impact from current choices (i.e., the selection of $x_{m}$), thereby allowing for a more informed and potentially more accurate selection of candidates.
Algorithm 2 Search-based Decoding
1:
2:$V$triangleright Vocabulary
3:$P_{tar},P_{para}$triangleright Target and paraphrase models
4:$k$triangleright Number of candidate sequences to retain
5:
6:$x_{best}$triangleright Best generated sequence
7:$X_{0}\leftarrow\{\emptyset\}$triangleright Initialize with empty sequence
8:while not $end$ do
9:     $m\leftarrow$ current decoding step
10:     $X^{\prime}_{m}\leftarrow\emptyset$
11:     for each $\mathrm{x}_{<m}\in X_{m-1}$ do
12:         for each $x_{m}\in V$ do
13:              Append $x_{<m}\circ x_{m}$ to $X^{\prime}_{m}$
14:         end for
15:     end for
16:     Use $P_{para}$ for $k$ candidates for each $\mathrm{x}_{<m}$
17:     $X_{m}\leftarrow\underset{(\mathrm{x}_{<m},x_{m})\in X^{\prime}_{m}}{\mathrm{\arg\,topk}}\{f(\mathrm{x}_{<m},x_{m})\}$triangleright Select $k$ candidates with lowest perplexity
18:     $end\leftarrow$ If any sequence ends with $EOS$
19:end while
20:$x_{best}\leftarrow$ select the sequence from $X_{m}$ with the lowest perplexity by $P_{tar}$
21:return $x_{best}$
3 Experiment
In this section, we demonstrate two distinct experimental settings for evaluation. In Section 3.1, we examine MonoPara’s effectiveness on prompt refinement for eliciting better task performance. In Section 3.2, we evaluate MonoPara’s effect on enhancing model robustness and generalization under instruction perturbations.
3.1 Task I: Prompt Refinement
Task Description
To inspect if monotonic paraphrasing can refine a prompt and elicit better performance on prompting an LLM as designed, we use Super-NaturalInstruction (Sup-NatInst for short, Wang et al., 2022). This benchmark consists of 1,616 diverse NLP tasks and their expert-written prompts for evaluating the zero-shot generalizability of LLMs on a variety of NLP tasks. In Sup-NatInst, each task is paired up with an instruction that consists of the task definition for mapping an input text to a task output and several examples for demonstrating the desired or undesired output. To make our evaluation more challenging, we only use the task definition as a prompt. Since Sup-NatInst provides only one description of definition for each task, we use GPT4 Achiam et al. (2023) to generate $4$ more Sup-NatInst-like task descriptions for each involved task, which results in $5$ prompts for each task in total.
Datasets
Following Gonen et al. (2023), we choose $7$ classification tasks from Huggingface Datasets  [^2] https://huggingface.co/docs/datasets/index to have a set of diverse tasks: (i) GLUE Cola Warstadt et al. (2019) for grammatical acceptability discrimination; (ii) Newspop Moniz and Torgo (2018) for news classification; (iii)AG News Zhang et al. (2015) for news classification; (iv) IMDB Maas et al. (2011) for movie review sentiment analysis; (v) DBpedia Lehmann et al. (2015) for topic classification; (vi) Emotion Saravia et al. (2018) for tweet emotion classification; and (vii) Tweet Offensive Barbieri et al. (2020) for offensive tweet discrimination. We sample 1,000 random examples from each dataset for prompt evaluation. For these tasks, the prompt follows the input, and at the end of each prompt, we add the choices of classes following Gonen et al. (2023): “Choices: X, Y, Z. Answer: ” as it helps with accuracy. Defining the label space likely helps in the zero-shot setting because there are no previous demonstrations from which the model can learn the possible classes.
Evaluation Metrics
To inspect whether monotonic paraphrasing can refine a prompt into its low-perplexity counterpart and further enhance the generalization of LLM prompting, we use PPL (perplexity), Acc (accuracy), and BS (BERTScore, Zhang* et al., 2020) for evaluation. Perplexity. Following Gonen et al. (2023), we define the perplexity of the prompt as the perplexity of the full prompt sequence, including the input itself and the choices of labels. To avoid noise when computing perplexity, the prompts are instantiated with 1,000 random examples of the dataset (we keep the same selected examples for performance evaluation), and the perplexity is averaged across all instantiated prompts. BERTScore. As a paraphrasing scheme, MonoPara is also evaluated by its performance in semantic alignment with the source sentence. BERTScore computes a similarity score for each token in the paraphrase sentence with each token in the source sentence, which is evaluated based on pre-trained BERT Devlin et al. (2019) contextual embeddings. Accuracy. To compute accuracy, for each test sample, we obtain the model’s predicted probability of all classes and choose the highest ranking class as the prediction of the model. We use this metric over exact match for its stableness since the chances the model will generate an exact match of the label are relatively low and might appear in various formats. The accuracy is calculated over the 1,000 samples for each task.
Methods Metric GLUE Cola Newspop AG News IMDB DBpedia Emotion Tweet Offensive
Mistral 7B
Acc 61.76 ± 2.93 85.28 ± 3.45 67.68 ± 3.47 84.12 ± 4.89 70.82 ± 2.27 44.32 ± 1.57 67.04 ± 0.95
Ori. PPL 15.72 ± 4.60 12.60 ± 2.40 12.74 ± 1.07 16.41 ± 0.80 8.10 ± 0.22 16.48 ± 2.14 19.25 ± 3.43
Acc 39.48 ± 8.19 76.72 ± 16.54 71.60 ± 2.70 86.60 ± 4.86 70.22 ± 5.71 43.65 ± 3.41 65.04 ± 1.46
PPL 78.01 ± 11.22 49.37 ± 7.81 28.37 ± 2.58 22.27 ± 0.98 28.34 ± 1.36 46.53 ± 5.30 53.65 ± 1.92
Greedy BS $8$ 0.74 87.47 ± 0.84 86.32 ± 2.08 90.23 ± 0.99 86.96 ± 0.93 90.08 ± 0.35 89.93 ± 0.67
Acc 63.96 ± 3.88 88.96 ± 3.04 72.20 ± 3.87 87.58 ± 3.27 72.07 ± 3.54 47.76 ± 3.72 67.00 ± 1.21
PPL 14.89 ± 2.75 8.33 ± 1.30 10.07 ± 2.93 13.05 ± 1.86 8.85 ± 0.88 13.67 ± 1.87 14.11 ± 3.53
Mono-E BS 93.03 ± 2.15 92.85 ± 2.35 88.70 ± 2.35 89.35 ± 2.60 93.21 ± 1.16 92.74 ± 2.75 91.73 ± 1.32
Acc 62.31 ± 6.17 86.34 ± 5.16 70.80 ± 3.44 83.48 ± 2.87 72.46 ± 3.00 44.26 ± 3.05 65.66 ± 2.43
PPL 15.85 ± 4.88 11.95 ± 2.09 9.89 ± 1.31 16.38 ± 1.63 8.63 ± 0.41 15.03 ± 1.65 18.35 ± 4.07
Mono-S BS 92.47 ± 2.00 93.78 ± 2.68 91.57 ± 2.23 89.66 ± 3.25 95.87 ± 1.20 93.42 ± 1.02 91.62 ± 2.48
Starling 7B
Acc 55.18 ± 12.43 93.78 ± 1.18 88.44 ± 1.18 96.46 ± 0.17 94.06 ± 0.37 55.48 ± 0.91 64.62 ± 0.18
Ori. PPL 16.38 ± 4.18 12.77 ± 2.59 12.10 ± 0.91 15.56 ± 0.57 8.11 ± 0.22 17.90 ± 2.82 20.41 ± 3.90
Acc 40.38 ± 15.49 82.67 ± 19.63 86.39 ± 2.92 96.66 ± 0.34 93.10 ± 1.79 51.10 ± 10.18 64.44 ± 0.08
PPL 70.50 ± 16.38 54.41 ± 8.80 30.67 ± 2.81 21.78 ± 1.37 26.56 ± 1.39 58.87 ± 4.44 11.71 ± 0.86
Greedy BS 90.35 ± 0.67 88.35 ± 0.51 88.46 ± 0.98 90.04 ± 1.04 87.06 ± 0.79 90.38 ± 0.45 79.93 ± 11.52
Acc 69.62 ± 4.12 95.15 ± 0.59 85.40 ± 4.22 96.22 ± 0.47 94.05 ± 0.42 56.10 ± 1.08 64.45 ± 0.09
PPL 8.56 ± 1.00 10.12 ± 1.02 8.52 ± 0.76 12.60 ± 1.57 9.58 ± 0.86 9.23 ± 0.97 11.71 ± 0.86
Mono-E BS 93.00 ± 0.85 94.19 ± 0.94 91.06 ± 1.51 92.85 ± 1.39 93.23 ± 0.50 93.28 ± 0.26 93.69 ± 1.52
Acc 50.96 ± 11.52 88.20 ± 3.74 83.61± 6.30 92.96 ± 4.78 92.82 ± 2.17 54.90 ± 2.32 66.80 ± 1.22
PPL 14.05 ± 3.76 12.55 ± 2.07 10.61 ± 1.06 16.44 ± 0.92 8.39 ± 0.53 16.28 ± 2.18 20.17 ± 3.17
Mono-S BS 92.69 ± 0.99 95.83 ± 1.66 93.92 ± 2.23 93.01 ± 1.25 96.99 ± 0.70 94.78 ± 1.25 94.45 ± 1.43
Table 1: Results of Task I (Prompt Refinement). Ori. refers to the original prompts. Greedy refers to paraphrasing the original prompts without perplexity constraint. Mono-E refers to ensemble-based MonoPara. Mono-S refers to search-based MonoPara. The best accuracy for each task is in bold, while the lowest perplexity for each task is in blue.
Models
We study two auto-regressive models, Mistral-7B Jiang et al. (2023)and Starling-7B Zhu et al. (2023a), for their significant ability in instruction following and paraphrasing.
Results
As shown in Table 1, the two variants of MonoPara successfully refines prompts to those with perplexities lower than the original ones, thereby enhancing model performance on 10 out of 14 scenarios (2 models, each on 7 datasets). Specifically, Mono-E (ensembled-based MonoPara) improves the accuracy of Mistral-7B by an average of 2.64% and that of Starling-7B by an average of 1.85% on seven different datasets. In contrast, directly paraphrasing the prompts without perplexity constraint results in prompts with higher perplexities than the original ones in 13 out of 14 scenarios, most of which lead to worse task performance. This highlights the importance of incorporating perplexity constraint when paraphrasing prompts. Moreover, MonoPara achieves superior performance in terms of both average prompt perplexity and BertScore, indicating the synergistic role played by both the target model and the paraphrase model in prompt refinement. During decoding, the target model prioritizes achieving low perplexity, while the paraphrase model preserves the original prompt intent, resulting in a high BertScore.
3.2 Task II: Robustness Evaluation on Instruction Perturbation
We further evaluate MonoPara’s effectiveness in enhancing model robustness against various instruction perturbations at character, word, sentence, and semantic levels using PromptBench Zhu et al. (2023b).
Task Description
PromptBench introduces perturbation to task instructions of a diverse set of tasks, such as sentiment analysis, grammar correctness, duplicate sentence detection, and natural language inference. It includes perturbations at character-level using DeepWordBug Gao et al. (2018) to introduce typos, word-level using TextFooler Jin et al. (2020) to replace words with contextually similar words, and semantic-level by using prompts following the linguistic behavior of different languages.
Datasets
We select $4$ GLUE tasks Wang et al. (2018) from Huggingface Datasets for evaluation: (i) Cola, which is also used for Task I; (ii) SST-2 Socher et al. (2013) for two-way sentiment classification; (iii) MNLI Wang et al. (2018) for multi-genre inference identification; and (iv) MRPC Dolan and Brockett (2005) for paraphrase identification. Similar to the setting of Task I, the instruction follows the input, and at the end of each instruction, the choices of labels are appended: “Choices: X, Y, Z. Answer: .”
Model
For this task, we use Alpaca Taori et al. (2023), a model instruction-tuned from the LLaMA model Touvron et al. (2023a) on the 52k self-instruct dataset.
Results
As shown in Figure 3, we compare the performance of original prompts, prompts paraphrased without perplexity constraint, and our method Mono-E. Across all kinds of instruction perturbations, our method achieves consistent accuracy improvement by up to 10.13%. This demonstrates that MonoPara not only refines prompts for improved task performance but also enhances model robustness against various instruction perturbations.
3.3 Effect of Coefficient
We further analyze the effect of coefficient $\alpha$ between the target model and the paraphrase model in MonoPara. Table 2 presents a case study of $\alpha$ on the ensemble-based decoding variant (Mono-E) with a prompt under character-level perturbation. Obviously, a larger value of $\alpha$ amplifies the influence of the target model, resulting in a stronger perplexity constraint and monotonically lower perplexity. Consequently, the impact of the paraphrase model is diminished, leading to a lower BertScore. By adjusting $\alpha$, the user can find a balanced point to achieve better prompts that strike a balance between semantic fidelity to the original prompt intent and linguistic familiarity to the target model.
Content BertScore PPL
Ori Instruction
Read the provided excerpt and choose between
‘positive’ and ‘negative’ to describe its sentiment Iyk1lJt8yw:
- 78.12
Greedy Para Determine the sentiment of the given excerpt as either positive or negative 81.26 44.29
Mono $\alpha=0.2$ Determine the sentiment of the given text by selecting either ‘positive’ or ‘negative’. 86.18 24.90
$\alpha=0.5$ Determine the sentiment of the given text by classifying it as positive or negative. 80.22 16.27
[…]
Table 2: Effect of coefficient $\alpha$ on Mono-E with Mistral-7B model.
Figure 3: Model’s average accuracy across $4$ GLUE datasets, with each dataset having six instructions with perturbation added at character, word, and semantic levels. Mono-E has consistent improvement in accuracy across all types of perturbation compared to greedy paraphrasing.
4 Related Work
Paraphrasing. Paraphrase generation has been a longstanding task in NLP. Since the era of deep learning, studies have adopted the universal paradigm of training sequence-to-sequence generation models based on abundant learning resources Nema et al. (2017); Gupta et al. (2018); Li et al. (2018). While more recent works have leveraged more robust pre-trained language models that captured rich supervision signals from various sequence-to-sequence generation tasks Raffel et al. (2020); Lewis et al. (2020) and achieving even more robust performance of paraphrasing with relatively limited end-task supervision Sun et al. (2021); Meng et al. (2021); Bui et al. (2021). In addition to these advancements, a series of works have added auxiliary supervision signals such as syntactical guidance, lexical regularization, and cross-lingual supervision, to further improve the performance and controllability of paraphrasing in aspects such as syntactic structures, stylistic specification, or textual simplicity Iyyer et al. (2018); Li et al. (2019); Chen et al. (2019); Kumar et al. (2020); Goyal and Durrett (2020); Hosking and Lapata (2021); Chowdhury et al. (2022).
Different from existing paraphrasing techniques, this work conducts constrained decoding methods that seek to monotonically decrease the perplexity in the process of paraphrasing. This innovative approach serves as a general approach to rewrite and refine prompts or task instructions towards their more familiar counterparts for LMs, seeking to enhance the generalization of LM prompting without the need of any training effort.
Prompt Refinement. Prompt refinement has been actively explored in recent years with the aim of selecting, retrieving, or even generating prompts that lead to improved zero-shot performance of LM prompting. When an end-task objective is known, prior works on prompt refinement approaches often set the end-task performance as the objective, and leverage approaches such as gradient-based search, similar to continuous prompts but with projections onto a discrete vocabulary Shin et al. (2020). Other works have explored edit-based enumeration Prasad et al. (2023), reinforcement learning Deng et al. (2022), and large language model continuation and filtering Zhou et al. (2022). As for general purpose LLMs, a dedicated prompt refinement approach may not assume the pre-existance of any specific end tasks. Hence, more recent works Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023) explored the heuristic criterion based on the familiarity of the language model with the language contained in the prompt, as measured by its perplexity. Lower perplexity prompts are preferred as they are expected to perform better across a wide range of tasks.
While the last line of research has relieved the generic principle for improving zero-shot LLM prompting generalizability, these studies are largely limited to selecting among pre-existing prompts or multiple beam search samples. Our proposed monotonic paraphrasing approach, however, can proactively rewrite a prompt or task instruction into counterparts that better satisfy the above heuristic criterion, without the need of any pre-existing candidate prompts or any training effort.
Ensemble and Search-based Decoding. The proposed two decoding schemes of MonoPara are connected to both ensemble decoding and search-based decoding approaches in recent studies. Ensemble decoding has been previously proposed for purposes including debiasing Li et al. (2023), controllable generation Meng et al. (2022); Huang et al. (2023), and cross-modality data processing Liu et al. (2023). Search-based decoding, on the other hand, is proposed to efficiently extend the search space of generation samples for satisfying specific constraints or criteria Lu et al. (2022a); Wuebker et al. (2012); Lee and Berg-Kirkpatrick (2022); Xu et al. (2023). Representative approaches include look-ahead Lu et al. (2022a) and look-back Xu et al. (2023) ones that search within the beam space towards different directions.
The explored decoding schemes in MonoPara are inspired by both lines of studies on decoding approaches. Specifically, the first line is related to our design choice for composing the probabilities of both the paraphrasing and the target LMs in greedy generation. And our search-based decoding scheme is specifically inspired by the look-ahead decoding strategy in the second line of studies.
5 Conclusion
In this paper, we propose monotonic paraphrasing (MonoPara) for automatically and iteratively generating low-perplexity prompts for given LLMs, that can lead to better task performance. Following the high-level idea of paraphrasing prompts with perplexity constraints from the target LLM, we design ensemble-based and search-based decoding strategies for efficient prompt refinement. Experiments on prompt variations of unseen tasks and instructions demonstrate the effectiveness of MonoPara in reducing prompt perplexity while enhancing task performance and model robustness. Future work may apply MonoPara on other scenarios, such as searching stealthy prompts for red-teaming. Replacing perplexity with other criteria for different purposes of constrained generation are also meaningful research directions.
Limitations
The current investigation of MonoPara has the following limitations. First, using the target model as the paraphrase model is a natural technical choice. However, examining other paraphrasing models, especially those customized for paraphrasing, may lead to better performance. Second, we apply look-ahead decoding to improve search efficiency. However, exploring more advanced decoding methods, such as look-back decoding Xu et al. (2023), may also prove beneficial. Third, while we conduct experiments on a wide range of tasks, their original prompts and responses are relatively short. This actually limits the operational space of MonoPara. Experiments on tasks with longer inputs and outputs may provide additional evidence of MonoPara’s effectiveness.